# UCFE: A User-Centric Financial Expertise Benchmark for Large Language Models

### v2

@@ Abstract
This paper introduces the UCFE: User-Centric Financial Expertise benchmark, an innovative framework designed to evaluate the ability of large language models (LLMs) to handle complex real-world financial tasks. UCFE benchmark adopts a hybrid approach that combines human expert evaluations with dynamic, task-specific interactions to simulate the complexities of evolving financial scenarios. Firstly, we conducted a user study involving 804 participants, collecting their feedback on financial tasks. Secondly, based on this feedback, we created our dataset that encompasses a wide range of user intents and interactions. This dataset serves as the foundation for benchmarking 12 LLM services using the LLM-as-Judge methodology. Our results show a significant alignment between benchmark scores and human preferences, with a Pearson correlation coefficient of 0.78, confirming the effectiveness of the UCFE dataset and our evaluation approach. UCFE benchmark not only reveals the potential of LLMs in the financial sector but also provides a robust framework for assessing their performance and user satisfaction.The benchmark dataset and evaluation code are available 111https://github.com/TobyYang7/UCFE-Benchmark.
@@
Yuzhe Yang1,††thanks: Equal contribution., Yifei Zhang2,11footnotemark: 1, Yan Hu1, Yilin Guo1, Ruoli Gan1, Yueru He3, Mingcong Lei1, Xiao Zhang3, Haining Wang2, Qianqian Xie3,22footnotemark: 2, Jimin Huang3, Honghai Yu2,, Benyou Wang1,††thanks: Corresponding author. 1The Chinese University of Hong Kong, Shenzhen, 2Nanjing University, 3The Fin AI yuzheyang@link.cuhk.edu.cn, yf_zhang@smail.nju.edu.cn,
@@ 1 Introduction
Recent advances in large language models (LLMs) have expanded their potential applications in finance Wu et al. (2023); Huang et al. (2023); Kim et al. (2024). Finance professionals are increasingly using LLMs to solve specialized financial tasks Li et al. (2023a); Xie et al. (2024); Yang et al. (2023); Zhang et al. (2023b). Financial tasks are inherently complex, involving specialized context, financial terminologies, legal intricacies, and dynamic markets that involve information with high noise-to-signal ratioPagano (1993); Mullainathan and Spiess (2017); Li et al. (2018), which adds significant challenges for LLMs to address. Accurate analysis of financial information is crucial, as even minor ignorance in a signal or market information can lead to substantial financial losses Tversky and Kahneman (1981); Thaler (2008); Mohamed et al. (2024).
To be effective, LLMs need to swiftly adapt to fiscal policy changes, market fluctuation, extreme events, and global factors, identifying key signals within real-time data to manage volatility and mitigate risks Gueta et al. (2024); Yadav et al. (2024). Financial markets can react instantly to news, making it crucial for LLMs to process information in near real-time by rapidly consolidating unstructured, real-time data from multiple sources Nguyen and Tulabandhula (2023); Tong et al. (2024). Despite LLMs’ improving accuracy on tasks like sentiment analysis, market prediction, and risk assessment Wimmer and Rekabsaz (2023); Lopez-Lira and Tang (2023); Rizinski et al. (2024), these models still face significant limitations, such as their reliance on static datasets and challenges in handling real-time data, which hinder their real-world applicability in dynamic financial contexts. Moreover, the evolving nature of financial regulation adds another layer of complexity, requiring LLMs to continuously update their knowledge to remain compliant and useful Yao et al. (2024); He et al. (2024). These limitations highlight the need for a more dynamic evaluation framework that assesses LLMs’ performance under real-time, evolving financial conditions, ensuring they can handle not only static tasks but also the unpredictable nature of real-world financial environments.
@@ 2.1 Financial Benchmark
FLARE Xie et al. (2023) evaluates models on five financial tasks 222sentiment analysis, news headline classification, named entity recognition, question answering, and stock movement prediction. Existing benchmarks Zhang et al. (2023a); Li et al. (2023b); Yuan et al. (2024) primarily use multiple-choice questions to assess models’ domain knowledge, with questions sourced from real-world financial documents and publicly available financial reports or websites, covering a wide range of topics such as finance, economy, accounting, and certification. MMMU and MMMU-PRO Yue et al. (2024a, b) extend beyond traditional financial NLP tasks by incorporating multimodal inputs to better evaluate models in more complex financial tasks. Although these benchmarks have advanced the evaluation of financial language models, they predominantly consider structured NLP tasks with deterministic answers and rely heavily on multiple-choice questions or tasks with specific answers, such as sentiment analysis and named entity recognition. This limits their ability to assess generative capabilities, which are essential for simulating real-world financial applications Krause (2023); Koa et al. (2024).
@@ 3 Background
Recent advancements in LLMs have demonstrated significant potential in addressing complex financial tasks. Numerous organizations are now actively training their own LLMs, aiming to enhance their performance by incorporating extensive domain-specific knowledge. For instance, FinGPT Yang et al. (2023), which applies supervised fine-tuning to the LLaMA model, has shown notable improvements in financial tasks. While models like FinLLaMA Xie et al. (2024), through continued pretraining, have further advanced LLM performance across various metrics. These developments highlight the growing demand and potential for LLMs in the financial domain, both in academia and industry.
@@
As discussed in Section 2.1, existing benchmarks largely focus on technical metrics such as accuracy and efficiency, often ignoring these broader challenges. By emphasizing only technical aspects, such benchmarks fail to address the real-world complexities of financial applications, where business rules, legal frameworks, and human judgment play crucial roles. This leaves non-technical aspects, such as human-AI interaction in finance, relatively under-explored. Human-AI interaction is critical in financial settings, as it affects decision-making, user trust, and the effective integration of AI systems into the financial workflow. Without considering these factors, current benchmarks offer an incomplete evaluation, limiting the practical relevance of LLMs for real-world financial applications. There is a pressing need for evaluation frameworks that not only assess technical performance but also account for the nuanced interplay between AI systems and financial professionals in complex environments.
In addition to these challenges, the rise of FinTech companies such as Robinhood 333https://robinhood.com/us/en/ has spurred increasing public interest in finance and trading. A growing number of individuals, many without formal financial education, are seeking accessible ways to manage their finances and participate in the market. For these users, LLMs have become a significant source of financial learning and advice. However, the accuracy, flexibility, and contextual understanding of LLMs are crucial to meet this emerging demand. Models must not only provide correct and relevant information but also adapt to the diverse financial needs of the general public, from basic education to advanced financial management strategies.
@@ 5 UCFE Benchmark
In this section, we provide an overview of the technical details and evaluation pipeline of the UCFE benchmark. As shown in Figure 4, the evaluation starts by selecting finance-specific tasks (introduced in Section 4.3), where the model acts as an AI assistant. GPT-4o444https://openai.com/index/gpt-4o-system-card/ simulates user interactions, generating dialogue data based on realistic behavior. Using LLMs to simulate user roles is common in recent research Inaba et al. (2024). To minimize model bias, we established evaluation criteria (detailed in Section 5.2). Model outputs are then compared in pairs, with Claude-3.5-Sonnet555https://www.anthropic.com/news/claude-3-5-sonnet as the evaluator, following the common practice LLm-as-judge framework for evaluation Liu et al. (2023). Each model’s performance is measured using Elo scores, which offer a relative comparison of capabilities. Finally, results are compared against human expert preferences to ensure evaluation robustness.
@@ 5.1 Evaluation Method
Each model starts with an Elo rating of , which is updated after every comparison task. For each task, dialogues generated by the target model and the base model are compared using specific prompts. A GPT-based model evaluates the comparison as a win, loss, or tie, and the Elo ratings are updated using the formula:
@@ 5.2 Experimental Settings
The backbone model of Tongyi-Finance-14B-Chat is Qwen-14B. Model Card: https://modelscope.cn/models/TongyiFinance/Tongyi-Finance-14B/summary
@@ 5.3 Overall Results
Table 5 shows the results across different tasks. Models like Tongyi-Finance-14B-Chat and CFGPT2-7B, which were continued pre-trained on financial corpus, exhibit significant improvements over their original backbone LLMs.
To ensure robustness, we used two additional evaluators, Gemini-1.5-pro666https://deepmind.google/technologies/gemini/pro/ and Deepseek-chat777https://www.deepseek.com/. As seen in Figure 5, the final results across target models show minimal variation, suggesting that the LLM-as-evaluator framework introduces slight bias. Additionally, to examine potential length bias Wei et al. (2024), we analyzed dialogue lengths and turn counts across models (Figure 6). The analysis shows no significant correlation, indicating no observable bias.
@@ 5.4 Human Preference Alignment
We invited 15 participants, including finance-related students and professionals, with each assigned 10 pre-existing results to conduct the human preference alignment phase. After manually reassessing the outputs, we updated the Elo scores and compared them with our model’s predictions. To quantify the similarity between the model’s results and human evaluations, we used the Pearson correlation coefficient:
@@ 5.5 Case Study
Previous research has demonstrated the significant influence of scaling laws on model performance Kaplan et al. (2020); Ruan et al. (2024). As shown in Figure 8 our results also show a similar trend. LLMs with larger parameters generally outperform smaller ones within the same series, also LLMs from the same backbone model have better results after being trained on financial corpus. However, Llama3.1 appears to be an outlier in this pattern. As shown in Figure 9, we illustrate a result of the Summarization task that highlights this phenomenon. In many of the data points, we observe that the 70B model tends to generate longer, more verbose outputs compared to smaller models. This aligns with the conclusions of Chiang and yi Lee (2024), where larger models are prone to over-reasoning, generating lengthy and unnecessary responses to questions. In our evaluation framework, where clear scoring criteria have been established, these redundant outputs significantly lower the model’s performance.
@@ Appendix A Questionnaire
The questionnaire of our survey is shown in Figure 10.
@@ B.1 Geographical Distribution
Figure 11 shows the geographical distribution of our previous survey. Among our 804 participants, 62.9% of them are from China, 35.9% from the USA, and 1.2% from other regions. This highlights the dominance of responses from China and the USA.
@@ B.2 Results of the survey
Figure 12 and Figure 13 report the primary financial information source and the results of whether users prefer generation answers or predefined options, which demonstrates the diversity of our benchmark contributors.
@@ Appendix C Human Expert Evaluation
Figure 15 illustrates the evaluation pipeline. Human experts first assess the existing results, and if the outcomes directly contradict their expertise, they are asked to make another selection. Additionally, we designed a UI interface, shown in Figure 16, to ensure that participants are unaware of the model names. After all the evaluations, we recalculated the Elo scores based on the samples evaluated by the human experts.
@@ Appendix D More Experiment Results
Figure 14 highlights each model’s comparative performance, showcasing strengths and weaknesses. And Table 6 presents the Elo scores across all tasks. The results show that LLMs trained on financial text data demonstrate outstanding performance across various tasks. However, due to limitations in the zero-shot task texts, the models’ performance in those tasks may be affected.
Figure 17 shows an example dialogue between User Simulator and Tongyi-Finance in the “Financial Knowledge Consulting” Task.
@@ E.1 Zero Shot Task
Figures [18, 19, 20] have shown all the prompts we used for testing and evaluation for Financial Text Summarization.
@@ E.2 Few Shot Task
Figures [21, 22, 23] have shown all the prompts we used for testing and evaluation for Asset Valuation Reporting.

### v3

@@ Abstract
This paper introduces the UCFE: User-Centric Financial Expertise benchmark, an innovative framework designed to evaluate the ability of large language models (LLMs) to handle complex real-world financial tasks. UCFE benchmark adopts a hybrid approach that combines human expert evaluations with dynamic, task-specific interactions to simulate the complexities of evolving financial scenarios. Firstly, we conducted a user study involving 804 participants, collecting their feedback on financial tasks. Secondly, based on this feedback, we created our dataset that encompasses a wide range of user intents and interactions. This dataset serves as the foundation for benchmarking 11 LLMs services using the LLM-as-Judge methodology. Our results show a significant alignment between benchmark scores and human preferences, with a Pearson correlation coefficient of 0.78, confirming the effectiveness of the UCFE dataset and our evaluation approach. UCFE benchmark not only reveals the potential of LLMs in the financial domain but also provides a robust framework for assessing their performance and user satisfaction.
@@
Yuzhe Yang1††thanks: Equal contribution: yuzheyang@link.cuhk.edu.cn, yf_zhang@smail.nju.edu.cn, huyan@cuhk.edu.cn, Yifei Zhang211footnotemark: 1, Yan Hu111footnotemark: 1, Yilin Guo1, Ruoli Gan1, Yueru He3, Mingcong Lei1, Xiao Zhang3, Haining Wang2, Qianqian Xie322footnotemark: 2, Jimin Huang3, Honghai Yu222footnotemark: 2, Benyou Wang1††thanks: Corresponding authors: qianqian.xie@thefin.ai, hhyu@nju.edu.cn, benyouwang@cuhk.edu.cn 1The Chinese University of Hong Kong, Shenzhen, 2Nanjing University, 3The Fin AI https://github.com/TobyYang7/UCFE-Benchmark
@@ 1 Introduction
Recent advances in large language models (LLMs) have expanded their potential applications in finance Wu et al. (2023); Huang et al. (2023); Kim et al. (2024). Finance professionals are increasingly using LLMs to solve specialized financial tasks Li et al. (2023a); Xie et al. (2024); Yang et al. (2023); Zhang et al. (2023b), including explorations into LLM-powered financial agents Li et al. (2024); Yang et al. (2025). Financial tasks are inherently complex, involving specialized context, financial terminologies, legal intricacies, and dynamic markets that involve information with high noise-to-signal ratio Pagano (1993); Mullainathan and Spiess (2017); Li et al. (2018), which adds significant challenges for LLMs to address. Accurate analysis of financial information is crucial, as even minor ignorance in a signal or market information can lead to substantial financial losses Tversky and Kahneman (1981); Thaler (2008); Mohamed et al. (2024).
To be effective, LLMs need to swiftly adapt to fiscal policy changes, market fluctuation, extreme events, and global factors, identifying key signals within real-time data to manage volatility and mitigate risks Gueta et al. (2024); Yadav et al. (2024). Financial markets can react instantly to news, making it crucial for LLMs to process information in near real-time by rapidly consolidating unstructured, real-time data from multiple sources Nguyen and Tulabandhula (2023); Tong et al. (2024). Despite LLMs’ improving accuracy on tasks like sentiment analysis, market prediction, and risk assessment Wimmer and Rekabsaz (2023); Lopez-Lira and Tang (2023); Rizinski et al. (2024), these models still face significant limitations, such as their reliance on static datasets and challenges in handling real-time data, which hinders their real-world applicability in dynamic financial contexts. Moreover, the evolving nature of financial regulation adds another layer of complexity, requiring LLMs to continuously update their knowledge to remain compliant and useful Yao et al. (2024); He et al. (2024). These limitations highlight the need for a more dynamic evaluation framework that assesses LLMs’ performance under real-time, evolving financial conditions, ensuring they can handle not only static tasks but also the unpredictable nature of real-world financial environments.
@@ 2.1 Financial Benchmark
FLARE Xie et al. (2023) evaluates models on five financial tasks 111sentiment analysis, news headline classification, named entity recognition, question answering, and stock movement prediction. Existing benchmarks Zhang et al. (2023a); Li et al. (2023b); Yuan et al. (2024) primarily use multiple-choice questions to assess models’ domain knowledge, with questions sourced from real-world financial documents and publicly available financial reports or websites, covering a wide range of topics such as finance, economy, accounting, and certification. MMMU and MMMU-PRO Yue et al. (2024a, b) extend beyond traditional financial NLP tasks by incorporating multimodal inputs to better evaluate models in more complex financial tasks. Although these benchmarks have advanced the evaluation of financial language models, they predominantly consider structured NLP tasks with deterministic answers and rely heavily on multiple-choice questions or tasks with specific answers, such as sentiment analysis and named entity recognition. This limits their ability to assess generative capabilities, which are essential for simulating real-world financial applications Krause (2023); Koa et al. (2024).
@@ 3 Background
Recent advancements in LLMs have demonstrated significant potential in addressing complex financial tasks. Numerous organizations are now actively training their own LLMs, aiming to enhance their performance by incorporating extensive domain-specific knowledge. For instance, FinGPT Yang et al. (2023), which applies supervised fine-tuning to the LLaMA model, has shown notable improvements in financial tasks. Through continued pretraining, models like FinLLaMA Xie et al. (2024) have further advanced LLM performance across various metrics. These developments highlight the growing demand and potential for LLMs in the financial domain, both in academia and industry.
@@
As discussed in Section 2.1, existing benchmarks largely focus on technical metrics such as accuracy and efficiency, often ignoring these broader challenges. By emphasizing only technical aspects, such benchmarks fail to address the real-world complexities of financial applications, where business rules, legal frameworks, and human judgment play crucial roles. This makes non-technical aspects, particularly human-AI interaction in finance, comparatively under-explored. Human-AI interaction is critical in financial settings, as it affects decision-making, user trust, and the effective integration of AI systems into the financial workflow. Without considering these factors, current benchmarks offer an incomplete evaluation, limiting the practical relevance of LLMs for real-world financial applications. There is a pressing need for evaluation frameworks that not only assess technical performance but also account for the nuanced interplay between AI systems and financial professionals in complex environments.
In addition to these challenges, the rise of FinTech companies such as Robinhood 222https://robinhood.com/us/en/ has spurred increasing public interest in finance and trading. A growing number of individuals, many without formal financial education, are seeking accessible ways to manage their finances and participate in the market. For these users, LLMs have become a significant source of financial learning and advice. However, the accuracy, flexibility, and contextual understanding of LLMs are crucial to meet this emerging demand. Models must not only provide correct and relevant information but also adapt to the diverse financial needs of the general public, from basic education to advanced financial management strategies.
@@ 5 UCFE Benchmark
In this section, we provide an overview of the technical details and evaluation pipeline of the UCFE benchmark. As shown in Figure 4, the evaluation starts by selecting finance-specific tasks (introduced in Section 4.3), where the model acts as an AI assistant. GPT-4o333https://openai.com/index/gpt-4o-system-card/ simulates user interactions, generating dialogue data based on realistic behavior. Using LLMs to simulate user roles is common in recent research Inaba et al. (2024). To minimize model bias, we established evaluation criteria (detailed in Section 5.2). Model outputs are then compared in pairs, with Claude-3.5-Sonnet444https://www.anthropic.com/news/claude-3-5-sonnet as the evaluator, following the common practice LLm-as-judge framework for evaluation Liu et al. (2023). Each model’s performance is measured using Elo scores, which offer a relative comparison of capabilities. Finally, results are compared against human expert preferences to ensure evaluation robustness.
@@ 5.1 Evaluation Method
Each model starts with an Elo rating of , which is updated after every comparison task. For each task, dialogues generated by the target model and the base model are compared using specific prompts. A Claude-based model evaluates the comparison as a win, loss, or tie, and the Elo ratings are updated using the formula:
@@ 5.2 Experimental Settings
The backbone model of Tongyi-Finance-14B-Chat is Qwen-14B.
@@ 5.3 Overall Results
Table 5 presents a comprehensive overview of model performance across the 17 distinct financial tasks within the UCFE benchmark. A key finding is the consistent outperformance of financially-specialized LLMs (Tongyi-Finance-14B-Chat, CFGPT2-7B, and Palmyra-Fin-70B-32k) compared to original backbone models ( Qwen series, InternLM series and Llama series). This performance gap is not uniform across all tasks, highlighting the varied strengths and weaknesses of different model architectures and training strategies when applied to the complexities of the financial domain. While general-purpose models demonstrate a baseline level of competence, their performance often lags significantly on tasks requiring in-depth financial knowledge, precise terminology, and adherence to regulatory constraints.
To ensure the robustness of our evaluation and mitigate potential biases inherent in using a single LLM as a judge (Claude-3.5-Sonnet), we conducted supplementary evaluations using two additional, independent LLM evaluators: Gemini-1.5-pro555https://deepmind.google/technologies/gemini/pro/ and Deepseek-chat666https://www.deepseek.com/. As illustrated in Figure 5, the final Elo scores across the target models show a high degree of consistency across all three evaluators, indicating that the relative rankings of the target models are largely insensitive to the choice of evaluator, suggesting that our findings are not an artifact of any particular evaluator’s biases or limitations.
Additionally, to examine potential length bias Wei et al. (2024), we analyzed dialogue lengths and turn counts across models (Figure 6). The analysis shows no significant correlation, indicating no observable bias.
@@ 5.4 Human Preference Alignment
To further validate that our results aligned with actual preferences of users, we also conducted a human preference evaluation involving 15 financial professionals and students, with each assigned 10 pre-existing results to conduct the human preference alignment phase. After manually reassessing the outputs, we updated the Elo scores and compared them with our model’s predictions. To quantify the similarity between the model’s results and human evaluations, we used the Pearson correlation coefficient:
@@ 5.5 Case Study
Previous research has demonstrated the significant influence of scaling laws on model performance Kaplan et al. (2020); Ruan et al. (2024). As shown in Figure 8, our results also show a similar trend. LLMs with larger parameters generally outperform smaller ones within the same series, also LLMs from the same backbone model have better results after being trained on financial corpus. However, Llama3.1 appears to be an outlier in this pattern. As shown in Figure 9, we illustrate a result of the Summarization task that highlights this phenomenon. In many of the data points, we observe that the 70B model tends to generate longer, more verbose outputs compared to smaller models. This aligns with the conclusions of Chiang and yi Lee (2024), where larger models are prone to over-reasoning, generating lengthy and unnecessary responses to questions. In our evaluation framework, where clear scoring criteria have been established, these redundant outputs significantly lower the model’s performance.
@@ 6 Conclusion
Acknowledgement
This work was supported by the Major Program of the National Fund of Philosophy and Social Science of China (No. 19ZDA105), the Shenzhen Science and Technology Program (JCYJ20220818103001002), the Shenzhen Doctoral Startup Funding (RCBS20221008093330065), the Tianyuan Fund for Mathematics of the National Natural Science Foundation of China (NSFC) (12326608), the Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (Grant No. ZDSYS20230626091302006), and the Shenzhen Stability Science Program 2023.
@@ Appendix A Questionnaire
The questionnaire of our survey is shown in Figure A-1.
@@ B.1 Geographical Distribution
Figure B-1 shows the geographical distribution of our previous survey. Among our 804 participants, 62.9% of them are from China, 35.9% from the USA, and 1.2% from other regions. This highlights the dominance of responses from China and the USA.
@@ B.2 Results of the survey
Figure B-2 and Figure B-3 report the primary financial information source and the results of whether users prefer generation answers or predefined options, which demonstrates the diversity of our benchmark contributors.
@@ Appendix C Human Expert Evaluation
Figure C-1 illustrates the evaluation pipeline. Human experts first assess the existing results, and if the outcomes directly contradict their expertise, they are asked to make another selection. Additionally, we designed a UI interface, shown in Figure C-2, to ensure that participants are unaware of the model names. After all the evaluations, we recalculated the Elo scores based on the samples evaluated by the human experts.
@@ Appendix D More Experiment Results
Figure B-4 highlights each model’s comparative performance, showcasing strengths and weaknesses. And Table D-1 presents the Elo scores across all tasks. The results show that LLMs trained on financial text data demonstrate outstanding performance across various tasks. However, due to limitations in the zero-shot task texts, the models’ performance in those tasks may be affected.
Figure C-3 shows an example dialogue between User Simulator and Tongyi-Finance in the “Financial Knowledge Consulting” Task.
@@ E.1 Zero Shot Task
Figures [E-1, E-2, E-3] have shown all the prompts we used for testing and evaluation for Financial Text Summarization.
@@ E.2 Few Shot Task
Figures [E-4, E-5, E-6] have shown all the prompts we used for testing and evaluation for Asset Valuation Reporting.